OpenStudy (anonymous):

help please x-3 over 4 =2 11 9 5 6

OpenStudy (anonymous):

algebra- (x-3)/4=2 multiply 4 to each side x-3=8 add 3 x=11 answer

OpenStudy (ivettef365):

so you have x - 3 ------ = 2 (what do you do first) 4

OpenStudy (anonymous):

11

OpenStudy (anonymous):

thank you all

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

yup